# MASt3R-Fusion: Integrating Feed-Forward Visual Model with IMU, GNSS for High-Functionality SLAM

### v2

@@ V Real-Time SLAM
V-A Isomorphic Group Transformation
To bridge the gap between Sim(3)-based visual alighment constraints and a more common metric-scale SE(3) factor graph, we introduce the isomorphic group transformation.
A Sim(3)-based similarity transformation can be described as
$\mathbf{S}\circ\mathbf{p}=\underbrace{\begin{bmatrix}s\mathbf{R}&\mathbf{t}\\ 0&1\end{bmatrix}}_{\mathbf{S}\in\mathrm{Sim}(3)}\begin{bmatrix}\mathbf{p}\\ 1\end{bmatrix}$, (20)
where $s\in\mathbb{R}^{+}$ denotes the scale, $\mathbf{R}\in\mathrm{SO}(3)$ the rotation, and $\mathbf{t}\in\mathbb{R}^{3}$ the translation.
@@ A Sim(3)-based similarity transformation can be described as
$\mathbf{T}\circ s\circ\mathbf{p}=\underbrace{\begin{bmatrix}\mathbf{R}&\mathbf{t}\\ 0&1\end{bmatrix}}_{\mathbf{T}\in SE(3)}\underbrace{\begin{bmatrix}s&0\\ 0&1\end{bmatrix}}_{\text{scaling}}\begin{bmatrix}\mathbf{p}\\ 1\end{bmatrix}$ (21)
@@
V-B Multi-Sensor Factor Graph
In the real-time SLAM stage, we maintain a sliding-window system state as follows
@@ V-B Multi-Sensor Factor Graph
where $N$ is the fixed size of the sliding window, and $K$ denotes the index of the latest keyframe, $\mathbf{T}_{i}\in\mathrm{SE}(3)$ is the camera pose (rotation and translation) of the $i$-th keyframe, $s_{i}\in\mathbb{R}$ is the scale parameter, $\mathbf{v}_{i}\in\mathbb{R}^{3}$ is the body velocity in the world frame, $\mathbf{b}_{i}=\left(\mathbf{b}_{i}^{g},\mathbf{b}_{i}^{a}\right)$ is the IMU bias vector, including gyroscope bias $\mathbf{b}_{i}^{g}\in\mathbb{R}^{3}$ and accelerometer bias $\mathbf{b}_{i}^{a}\in\mathbb{R}^{3}$. It should be noted that all states are maintained in double precision (float64).
For visual information, a quadric-form factor is utilized, following
@@
where $(\mathbf{H}_{ij}$, $\mathbf{v}_{ij})$ is the Hessian information of the Sim(3)-based visual constraint (Eq. (14)), $(\mathbf{H}_{v,ij}$, $\mathbf{v}_{v,ij})$ is the Hessian information after local-to-global transformation (through Jacobian $\mathbf{J}^{ij}_{(i,j)}$) and isomorphic group transformation, $l_{v}(\cdot)$ is the linear container that transform current state $\mathcal{X}$ to the Lie algebras at the linearization point in Eq. (11) .
When constructing visual information, we first obtain the relative Sim(3) transformation between projection frames using the maintained double-precision (float64) poses and scales. The Hessian information $(\mathbf{H}_{ij},\mathbf{v}_{ij})$ is then computed using Eq. (11) and (14) in single precision (float32), and subsequently transformed via a double-precision linear mapping (Eq. (30)) into constraints on the global parameters. By restricting single-precision GPU computation within the local scale, this strategy mitigates numerical instabilities in large-scale scenarios.
@@
For IMU information, the classic IMU pre-intergration[37] is used for state prediction and inter-frame constraints
$\begin{split}&\mathbf{r}_{b}\left(\mathcal{X}_{k},\mathcal{X}_{k+1}\right)=\\ &\left[\begin{matrix}{\mathbf{R}^{w}_{b_{k}}}^{\top}\left(\mathbf{p}^{w}_{b_{k+1}}\!-\!\mathbf{p}^{w}_{b_{k}}\!+\!\frac{1}{2}\mathbf{g}^{w}\Delta{t^{2}_{k}}\!-\!\mathbf{v}^{w}_{b_{k}}\Delta t_{k}\!\right)-\!\Delta\tilde{\mathbf{p}}^{b_{k}}_{b_{k+1}}\\ {\mathbf{R}^{w}_{b_{k}}}^{\top}\left(\mathbf{v}^{w}_{b_{k+1}}+\mathbf{g}^{w}\Delta t_{k}-\mathbf{v}^{w}_{b_{k}}\right)-\Delta\tilde{\mathbf{v}}^{b_{k}}_{b_{k+1}}\\ \text{Log}\left(\left({\mathbf{R}^{w}_{b_{k}}}\right)^{-1}\mathbf{R}^{w}_{b_{k+1}}{\left(\Delta\tilde{\mathbf{R}}^{b_{k}}_{b_{k+1}}\right)^{-1}}\right)\\ \mathbf{b}^{a}_{k+1}-\mathbf{b}^{a}_{k}\\ \mathbf{b}^{g}_{k+1}-\mathbf{b}^{g}_{k}\end{matrix}\right]\end{split}$ (31)
where $\Delta\tilde{\mathbf{p}}^{b_{k}}_{b_{k+1}}$, $\Delta\tilde{\mathbf{v}}^{b_{k}}_{b_{k+1}}$, $\Delta\tilde{\mathbf{R}}^{b_{k}}_{b_{k+1}}$ are the IMU preintegration terms[37], $\mathbf{g}^{w}$ is the gravity vector, $\Delta t_{k}$ is the time interval. Considering the difference between the camera pose and the IMU pose, the extrinsic parameters need to be taken into account, as shown in the following equation
@@
The fusion of visual and IMU information forms the fundamental mechanism of visual-inertial odometry (VIO), which can provide metric-scale information, low-drift pose estimation, and gravity awareness. To maximize the performance of this mechanism, probabilistic marginalization needs to be considered. To be specific, When the number of states within the sliding window exceeds $N$, a local factor graph associated with the oldest pose is constructed. The state at the oldest time step is then eliminated via the Schur complement[23], thereby yielding the marginalization information $(\mathbf{H}_{m},\mathbf{v}_{m})$. This leads to the following quadric-form factor
@@
Figure 6: Illustration of the sliding-window factor graph in the real-time SLAM stage. The Sim(3)-to-SE(3) isomorphic group transformation and IMU-camera extrinsic transformation are used to bridge the gap between visual constraints and inertial information. Probabilistic marginalization is employed for optimal information utilization with controlled problem scale. For clarity, only the adjacent visual constraints are shown.
Thus, the cost function corresponding to the real-time visual-inertial SLAM system is described as
@@
With the sliding-window setup, the scale of this optimization problem is kept small, and the system is free from the frequent re-computation of old projections. The computational load is mainly concentrated in the feed-forward model. The effective integration of V-I information allows the system to achieve high-precision pose/scale estimation using only local dense matching, enabling real-time operation on a 4080 laptop. Furthermore, the sliding-window design allows the system to handle arbitrarily long data sequences while requiring only 8 GB of GPU memory.
VI Global SLAM
@@ VI-A Loop Closure
First, as proposed in[6], we can use image feature tokens obtained from the feed-forward encoder to construct an image retrieval system
@@
Benefiting from the strong association capability of the feed-forward model, this system is highly effective to recognize subtle cues of a common scene, but also produces a considerable number of false positives. A further dense matching step can be used for verification of the loop closure, while this can introduce significant extra computational cost.
@@
Based on the odometry essence of VIO, we simplify the position estimation as a Markov process, leading to the following expression
$\Delta\hat{\mathbf{p}}^{w}_{p,q}=\sum_{p}^{q-1}{\Delta\hat{\mathbf{p}}^{w}_{i,i+1}}$ (37)
where $\Delta\hat{\mathbf{p}}^{w}_{p,q}$ is the relative translation estimation between timestamp $p$ and $q$. Considering the practical need of loop closure, we simplify the translation into 2D horizontal movement.
Considering the error of the odometry, we get
$\Delta\hat{\mathbf{p}}^{w}_{i,i+1}=\Delta\mathbf{p}^{w}_{i,i+1}+\bm{\epsilon}_{i,i+1},$ (38)
@@
By dividing the odometry uncertainty into “along” direction (corresponding to scale error) and “cross” direction (corresponding to heading error), we can estimate the translation uncertainty following
@@
where $d=\left\|\Delta\hat{\mathbf{p}}^{w}_{i,i+1}\right\|$ is the translation distance, $\mathbf{Q}$ is a direction-and-uncertainty-related ellipsoid, following
@@
$\begin{cases}\mathbf{P}_{\parallel}&=\mathbf{n}\mathbf{n}^{\top}\\ \mathbf{P}_{\perp}&=\mathbf{I}-\mathbf{n}\mathbf{n}^{\top}\\ \mathbf{n}&=\Delta\hat{\mathbf{p}}^{w}_{i,i+1}/\left\|\Delta\hat{\mathbf{p}}^{w}_{i,i+1}\right\|\end{cases}$ (41)
where $\sigma_{d}$ is the uncertainty factor related to the scale error of translation, $\sigma_{n}$ is the uncertainty factor related to the direction/heading.
Based on the above equations, we can efficiently compute the covariances of $\bm{\epsilon}_{i,i+1}$, and estimate the covariance of inter-frame translation $\Delta\hat{\mathbf{p}}^{w}_{p,q}$ by summing up. This operation is fast by utilizing vectorization. Afterwards, we turn the uncertainty of translation into the uncertainty of inter-frame distance
$\sigma_{p,q}=\sqrt{\frac{{\Delta\hat{\mathbf{p}}^{w}_{p,q}}^{\top}\mathbf{Q}_{p,q}\,{\Delta\hat{\mathbf{p}}^{w}_{p,q}}}{\left\|{\Delta\hat{\mathbf{p}}^{w}_{p,q}}\right\|^{2}}},$ (42)
where $\mathbf{Q}_{p,q}$ is the estimated covariance of translation $\Delta\hat{\mathbf{p}}^{w}_{p,q}$.
Equation (42) forms a 2D uncertainty map of inter-frame distances. It should be noted that, when predicting the co-visibility, we need to consider not only the spatial proximity of camera positions. Due to the presence of large-angle loop closures, co-visible frames may observe the same scene while exhibiting significant positional differences caused by varying viewpoints. Therefore, it is more important to focus on the proximity of observed points. To this end, we approximate the main observation point of a camera frame $i$ by introducing a point of interest, defined as
$\mathbf{p}_{o,i}=\left(\mathbf{T}_{i}\circ\begin{bmatrix}0\\ 0\\ L\end{bmatrix}\right)_{x,y}$ (43)
@@
Figure 7: Loop closure candidate filtering based on efficient positional uncertainty evaluation. The upper figure illustrates the set of loop closure candidates selected based on the similarity scores and temporal non-contiguity. After filtering with the distance uncertainty matrix $\sigma_{p,q}$, the lower figure presents the geometrically feasible loop closure candidates.
@@
$d\left(\mathbf{p}_{o,q},\mathbf{p}_{o,p}\right)<L$ (44)
The next task is to determine, based on positional uncertainty, whether there is a possibility that two frames satisfy condition (44). By approximating the uncertainty of the distance between points $\mathbf{p}_{o,p}$ and $\mathbf{p}_{o,q}$ to that of the distance between points $\mathbf{p}_{p}$ and $\mathbf{p}_{q}$, the probability distribution function corresponding to the above condition can be obtained
$P(d\left(\hat{\mathbf{p}}_{o,q},\hat{\mathbf{p}}_{o,p}\right)<L)=\Phi\left(\frac{L-d\left(\hat{\mathbf{p}}_{o,q},\hat{\mathbf{p}}_{o,p}\right)}{\sigma_{p,q}}\right)$ (45)
@@
On this basis, we take the following criterion to select possible loop closure candidates
$d\left(\hat{\mathbf{p}}_{o,q},\hat{\mathbf{p}}_{o,p}\right)<L+\sigma_{p,q}$ (46)
@@
A example of the candidate filtering is presented in Fig. 7.
VI-B Global Factor Graph Optimization
@@
For visual loop closure measurements, the relative pose estimate between loop closure frame $(i,j)$ can be directly obtained through the projection between two frames (Eq. (12)), which is $\tilde{\mathbf{S}}^{i}_{j,\mathrm{loop}}$. With the knowledge of $s_{i}$ obtained in the real-time SLAM stage, the SE(3) relative pose can be retrieved, which leads to a metric-scale relative pose contraint
@@
For more accurate constraints, we also extract the Hessian-form information $(\mathbf{H}_{ij,\mathrm{loop}},\mathbf{v}_{ij,\mathrm{loop}})$, $(\mathbf{H}_{ji,\mathrm{loop}},\mathbf{v}_{ji,\mathrm{loop}})$ during the estimation of $\tilde{\mathbf{S}}^{i}_{j,\mathrm{loop}}$.
@@
$\mathbf{r}_{g}(\mathcal{X}_{i})=\mathbf{T}^{n}_{w}\circ\mathbf{T}_{i}\circ\mathbf{T}^{b}_{c}\circ\mathbf{t}^{c}_{g}-\tilde{\mathbf{p}}^{n}_{g,i}$ (48)
where $\tilde{\mathbf{p}}^{n}_{g,i}$ is the measured position of the GNSS phase center in the navigation frame, $\mathbf{t}^{b}_{g}$ is the IMU-GNSS lever-arm, $\mathbf{T}^{n}_{w}$ is a fixed world-to-navigation transformation obtained through the initial alignment. To overcome the time asynchrony between GNSS measurements and keyframes, we introduce a temporary pre-integration term to bridge the GNSS measurement with the nearest keyframe timestamp. Thus, the combined cost function will be:
@@
During global optimization, due to the inconsistency between odometry poses and loop closure relative poses, directly using the Hessian-form loop information for optimization is ineffective (since the initial estimate deviates from the linearization point when estimating $\tilde{\mathbf{S}}^{i}_{j,\mathrm{loop}}$). Therefore, in the global pose graph, we first introduce loop closure constraints in the form of relative poses (47). The resulting optimization problem is formulated as follows:
@@
where $\mathcal{L^{\prime}}$ is the inlier loop closure set.
It is noted that during global optimization, the Hessian information $(\mathbf{H}_{ij},\mathbf{v}_{ij})$ with respect to the relative transformation $\mathbf{S}^{i}_{j}$ in Eq. (30) doesn’t need to be re-computed through direct projection Eq. (11), as the relative transformation is stable. Only the local-to-global linear mapping (Eq. (30)) needs to be iteratively performed, which makes the optimization highly efficient while keep the accurate visual information.
Through the stepwise optimization strategy and the precise utilization of measurement information, the global optimization problem can be solved efficiently while naturally detecting and mitigating outliers in GNSS and loop closure measurements in a probabilistic manner, thereby achieving nearly optimal estimation results. In contrast, traditional global fusion methods based solely on pose graphs discard the original IMU and visual information, causing inaccurate VIO pose estimates to significantly affect the global estimation, particularly with respect to scale uncertainty.
VII Experiments
In the experimental section, we first evaluate the performance of the visual-inertial integration on the public KITTI-360[38] and SubT-MRS[39] datasets, including both the real-time SLAM and global optimization stages. Subsequently, we assess the performance of the visual-inertial-GNSS fusion on our self-collected Wuhan urban dataset. Monocular camera information is used for all the tests.
VII-A KITTI-360 Dataset

### v3

@@ IV Real-Time SLAM
IV-A Isomorphic Group Transformation
To bridge the gap between Sim(3)-based visual alignment constraints and a more common metric-scale SE(3) factor graph, we introduce the isomorphic group transformation.
A Sim(3)-based similarity transformation can be described as:
$\mathbf{S}\circ\mathbf{t}=\underbrace{\begin{bmatrix}s\mathbf{R}&\mathbf{t}\\ 0&1\end{bmatrix}}_{\mathbf{S}\in\mathrm{Sim}(3)}\begin{bmatrix}\mathbf{t}\\ 1\end{bmatrix}$, (20)
@@ A Sim(3)-based similarity transformation can be described as:
$\mathbf{T}\circ s\circ\mathbf{t}=\underbrace{\begin{bmatrix}\mathbf{R}&\mathbf{t}\\ 0&1\end{bmatrix}}_{\mathbf{T}\in SE(3)}\underbrace{\begin{bmatrix}s&0\\ 0&1\end{bmatrix}}_{\text{scaling}}\begin{bmatrix}\mathbf{t}\\ 1\end{bmatrix}$ (21)
@@
IV-B Multi-Sensor Factor Graph
In the real-time SLAM stage, we maintain a sliding-window system state as follows:
@@ IV-B Multi-Sensor Factor Graph
where $N$ is the fixed size of the sliding window, and $K$ denotes the index of the latest keyframe, $\mathbf{T}_{i}\in\mathrm{SE}(3)$ is the camera pose (rotation and translation) of the $i$-th keyframe, $s_{i}\in\mathbb{R}^{+}$ is the scale parameter, $\mathbf{v}_{i}\in\mathbb{R}^{3}$ is the body velocity in the world frame, $\mathbf{b}_{i}=\left(\mathbf{b}_{i}^{g},\mathbf{b}_{i}^{a}\right)$ is the IMU bias vector, including gyroscope bias $\mathbf{b}_{i}^{g}\in\mathbb{R}^{3}$ and accelerometer bias $\mathbf{b}_{i}^{a}\in\mathbb{R}^{3}$. It should be noted that all states are maintained in double precision (float64).
For visual information, a quadric-form factor is utilized, following:
@@
where $(\mathbf{H}_{ij}$, $\mathbf{v}_{ij})$ is the Hessian information of the Sim(3)-based visual constraint (Eq. (14)), $(\mathbf{H}_{v,ij}$, $\mathbf{v}_{v,ij})$ is the Hessian information after relative-to-absolute transformation (through Jacobian $\mathbf{J}^{ij}_{(i,j)}$) and isomorphic group transformation (through $\bm{\Lambda}_{(i,j)}$), $l_{v}(\cdot)$ is the linear container that transform current state $\mathcal{X}$ to the Lie algebras at the linearization point in Eq. (11) .
When constructing visual information, we first obtain the relative Sim(3) transformation between projection frames using the maintained double-precision (float64) poses and scales. The local Hessian information $(\mathbf{H}_{ij},\mathbf{v}_{ij})$ is then computed using Eq. (11) and (14) in single precision (float32), and subsequently transformed via a double-precision linear mapping (Eq. (30)) into constraints on the global parameters. By restricting single-precision GPU computation within the local scale, this strategy mitigates numerical instabilities in large-scale scenarios.
@@
For IMU information, the classic IMU pre-intergration[39] is used for state prediction and inter-frame dynamic constraints:
$\begin{split}&\mathbf{r}_{b}\left(\mathcal{X}_{k},\mathcal{X}_{k+1}\right)=\\ &\left[\begin{matrix}{\mathbf{R}^{w}_{b_{k}}}^{\top}\left(\mathbf{t}^{w}_{b_{k+1}}\!-\!\mathbf{t}^{w}_{b_{k}}\!+\!\frac{1}{2}\mathbf{g}^{w}\Delta{t^{2}_{k}}\!-\!\mathbf{v}^{w}_{b_{k}}\Delta t_{k}\!\right)-\!\Delta\tilde{\mathbf{t}}^{b_{k}}_{b_{k+1}}\\ {\mathbf{R}^{w}_{b_{k}}}^{\top}\left(\mathbf{v}^{w}_{b_{k+1}}+\mathbf{g}^{w}\Delta t_{k}-\mathbf{v}^{w}_{b_{k}}\right)-\Delta\tilde{\mathbf{v}}^{b_{k}}_{b_{k+1}}\\ \text{Log}\left(\left({\mathbf{R}^{w}_{b_{k}}}\right)^{-1}\mathbf{R}^{w}_{b_{k+1}}{\left(\Delta\tilde{\mathbf{R}}^{b_{k}}_{b_{k+1}}\right)^{-1}}\right)\\ \mathbf{b}^{a}_{k+1}-\mathbf{b}^{a}_{k}\\ \mathbf{b}^{g}_{k+1}-\mathbf{b}^{g}_{k}\end{matrix}\right]\end{split}$ (31)
where $\Delta\tilde{\mathbf{t}}^{b_{k}}_{b_{k+1}}$, $\Delta\tilde{\mathbf{v}}^{b_{k}}_{b_{k+1}}$, $\Delta\tilde{\mathbf{R}}^{b_{k}}_{b_{k+1}}$ are the IMU preintegration terms[39], $\mathbf{g}^{w}$ is the gravity vector, $\Delta t_{k}$ is the time interval. Considering the difference between the camera pose and the IMU pose, the extrinsic parameters need to be taken into account, as shown in the following equation:
@@
The fusion of visual and IMU information forms the fundamental mechanism of visual-inertial odometry (VIO), which can provide metric-scale information, low-drift pose estimation, and gravity awareness. To maximize the performance of this mechanism, probabilistic marginalization needs to be considered. To be specific, When the number of states within the sliding window exceeds $N$, a local factor graph associated with the oldest pose is constructed. The state at the oldest time step is then eliminated via the Schur complement[7], thereby yielding the marginalization information $(\mathbf{H}_{m},\mathbf{v}_{m})$. This leads to the following quadric-form factor:
@@
Figure 6: Illustration of the sliding-window factor graph in the real-time SLAM stage. The Sim(3)-to-SE(3) isomorphic group transformation and IMU-camera extrinsic transformation are used to bridge the gap between visual constraints and inertial information. Probabilistic marginalization is employed for optimal information utilization with controlled problem scale. For clarity, only the adjacent visual constraints are depicted.
Thus, the cost function corresponding to the real-time visual-inertial SLAM system is described as:
@@
as is illustrated in Fig. 6.
With the sliding-window setup, the scale of this optimization problem is kept small, and the system is free from the frequent re-computation of older projections. In this case, the computational load is mainly concentrated in the feed-forward model. The effective integration of V-I information allows the system to achieve high-precision pose/scale estimation using only local dense matching, enabling real-time operation on a laptop with RTX 4080 Mobile GPU. Furthermore, the sliding-window design allows the system to handle arbitrarily long data sequences while requiring only 8 GB of GPU memory.
V Global SLAM
@@ V-A Loop Closure
First, as proposed in[5], we can use image feature tokens obtained from the feed-forward encoder to construct an image retrieval system:
@@
Benefiting from the strong association capability of the feed-forward model, this system is highly effective to recognize subtle cues of a common scene, but also produces a considerable number of false positives. A further dense matching step (see Sect. 4.1) can be used for verification of the loop closure, while this can introduce significant extra computational cost.
@@
Based on the odometry essence of VIO, we simplify the position estimation as a Markov process, leading to the following expression:
$\Delta\hat{\mathbf{t}}^{w}_{p,q}=\sum_{p}^{q-1}{\Delta\hat{\mathbf{t}}^{w}_{i,i+1}}$ (37)
where $\Delta\hat{\mathbf{t}}^{w}_{p,q}$ is the relative translation estimation between timestamp $p$ and $q$. Considering the practical need of loop closure, we simplify the translation into 2D horizontal movement.
Considering the error of the odometry, we get:
$\Delta\hat{\mathbf{t}}^{w}_{i,i+1}=\Delta\mathbf{t}^{w}_{i,i+1}+\bm{\epsilon}_{i,i+1},$ (38)
@@
By dividing the odometry uncertainty into “along” direction (corresponding to scale error) and “cross” direction (corresponding to heading error), we can estimate the translation uncertainty following:
@@
where $d=\left\|\Delta\hat{\mathbf{t}}^{w}_{i,i+1}\right\|$ is the translation distance, $\mathbf{Q}$ is a direction-and-uncertainty-related ellipsoid, following:
@@
$\begin{cases}\mathbf{P}_{\parallel}&=\mathbf{n}\mathbf{n}^{\top}\\ \mathbf{P}_{\perp}&=\mathbf{I}-\mathbf{n}\mathbf{n}^{\top}\\ \mathbf{n}&=\Delta\hat{\mathbf{t}}^{w}_{i,i+1}/\left\|\Delta\hat{\mathbf{t}}^{w}_{i,i+1}\right\|\end{cases}$ (41)
where $\sigma_{d}$ is the uncertainty factor related to the scale error of translation, $\sigma_{n}$ is the uncertainty factor related to the direction/heading. These two factors are preset parameters considering the normal VIO performance.
Based on the above equations, we can efficiently compute the covariances of $\bm{\epsilon}_{i,i+1}$, and estimate the covariance of inter-frame translation $\Delta\hat{\mathbf{t}}^{w}_{p,q}$ by summing up. This operation is fast by utilizing vectorization. Afterwards, we turn the uncertainty of translation into the uncertainty of inter-frame distance:
$\sigma_{p,q}=\sqrt{\frac{{\Delta\hat{\mathbf{t}}^{w}_{p,q}}^{\top}\mathbf{Q}_{p,q}\,{\Delta\hat{\mathbf{t}}^{w}_{p,q}}}{\left\|{\Delta\hat{\mathbf{t}}^{w}_{p,q}}\right\|^{2}}},$ (42)
where $\mathbf{Q}_{p,q}$ is the estimated covariance of translation $\Delta\hat{\mathbf{t}}^{w}_{p,q}$.
Equation (42) forms a 2D uncertainty map of inter-frame distances. It should be noted that, when predicting the co-visibility, we need to consider not only the spatial proximity of camera positions. Due to the presence of large-angle loop closures, co-visible frames may observe the same scene while exhibiting significant positional differences caused by varying viewpoints. Therefore, it is more important to focus on the proximity of observed points. To this end, we approximate the main observation point of a camera frame $i$ by introducing a point of interest, defined as:
$\mathbf{t}_{\bar{i}}=\left(\mathbf{T}_{i}\circ\begin{bmatrix}0\\ 0\\ L\end{bmatrix}\right)_{x,y}$ (43)
@@
Figure 7: Loop closure candidate filtering based on efficient positional uncertainty evaluation. The upper figure illustrates the set of loop closure candidates selected based on the similarity scores and temporal non-contiguity. After filtering with the distance uncertainty matrix $\sigma_{p,q}$ (Eq. 42), the lower figure presents the geometrically feasible loop closure candidates.
@@
$d\left(\mathbf{t}_{\bar{q}},\mathbf{t}_{\bar{p}}\right)<L$ (44)
where $d(\cdot,\cdot)$ is the Euclidean distance.
The next task is to determine, based on positional uncertainty, whether there is a possibility that two frames satisfy condition (44). By approximating the uncertainty of the distance between points $\mathbf{t}_{\bar{p}}$ and $\mathbf{t}_{\bar{q}}$ to that of the distance between points $\mathbf{t}_{p}$ and $\mathbf{t}_{q}$, the probability distribution function corresponding to the above condition can be obtained:
$P(d\left(\hat{\mathbf{t}}_{\bar{q}},\hat{\mathbf{t}}_{\bar{p}}\right)<L)=\Phi\left(\frac{L-d\left(\hat{\mathbf{t}}_{\bar{q}},\hat{\mathbf{t}}_{\bar{p}}\right)}{\sigma_{p,q}}\right)$ (45)
@@
On this basis, we take the following criterion to select possible loop closure candidates:
$d\left(\hat{\mathbf{t}}_{\bar{q}},\hat{\mathbf{t}}_{\bar{p}}\right)<L+\sigma_{p,q}$ (46)
@@
Examples of the above-described candidate filtering are presented in Fig. 7.
V-B Global Factor Graph Optimization
@@
For visual loop closure measurements, the relative pose estimate between loop closure frame $(i,j)$ can be directly obtained through the projection between two frames (Eq. (12)), which is $\tilde{\mathbf{S}}^{i}_{j,\mathrm{loop}}$. With the knowledge of $s_{i}$ obtained in the real-time SLAM stage, the SE(3) relative pose can be retrieved, which leads to a metric-scale relative pose contraint:
@@
For more accurate constraints, we also store the Hessian-form information $(\mathbf{H}_{ij,\mathrm{loop}},\mathbf{v}_{ij,\mathrm{loop}})$, $(\mathbf{H}_{ji,\mathrm{loop}},\mathbf{v}_{ji,\mathrm{loop}})$ during the estimation of $\tilde{\mathbf{S}}^{i}_{j,\mathrm{loop}}$.
@@
$\mathbf{r}_{g}(\mathcal{X}_{i})=\mathbf{T}^{n}_{w}\circ\mathbf{T}_{i}\circ\mathbf{T}^{b}_{c}\circ\mathbf{t}^{c}_{g}-\tilde{\mathbf{t}}^{n}_{g,i}$ (48)
where $\tilde{\mathbf{t}}^{n}_{g,i}$ is the measured position of the GNSS phase center in the navigation frame, $\mathbf{t}^{b}_{g}$ is the IMU-GNSS lever-arm, $\mathbf{T}^{n}_{w}$ is a fixed world-to-navigation transformation obtained through the initial alignment. To overcome the time asynchrony between GNSS measurements and keyframes, we introduce a temporary pre-integration term to bridge the GNSS measurement with the nearest keyframe timestamp. Thus, the combined cost function will be:
@@
During global optimization, due to the inconsistency between odometry poses and loop closure relative poses, directly using the Hessian-form loop information for optimization is instable (since the initial estimate deviates from the linearization point when estimating $\tilde{\mathbf{S}}^{i}_{j,\mathrm{loop}}$). Therefore, in the global pose graph, we first introduce loop closure constraints in the form of relative poses (47). The resulting optimization problem is formulated as follows:
@@
where $\mathcal{L^{\prime}}$ is the inlier loop closure set, the Hessian-form loop closure constraints are derived from $(\mathbf{H}_{ij,\mathrm{loop}},\mathbf{v}_{ij,\mathrm{loop}})$ and $(\mathbf{H}_{ji,\mathrm{loop}},\mathbf{v}_{ji,\mathrm{loop}})$.
It is noted that during global optimization, the Hessian information $(\mathbf{H}_{ij},\mathbf{v}_{ij})$ with respect to the relative transformation $\mathbf{S}^{i}_{j}$ in Eq. (30) doesn’t need to be re-computed through performing the dense projection Eq. (11), as the relative transformation is stable. Only the relative-to-absolute linear mapping (Eq. (30)) needs to be iteratively performed, which makes the optimization highly efficient while keep the accuracy of the visual information.
Through the stepwise optimization strategy and the precise utilization of measurement information, the global optimization problem can be solved efficiently while naturally mitigating outliers in GNSS and loop closure measurements in a probabilistic manner, thereby achieving nearly optimal estimation results. In contrast, traditional global fusion methods based solely on pose graphs discard the original IMU and visual information, causing inaccurate VIO pose estimates to significantly affect the global estimation, particularly with respect to the scale uncertainty.
VI Experiments
In the experimental section, we first evaluate the visual-inertial integration performance of the system on the public KITTI-360[40] and SubT-MRS[41] datasets, including both the real-time SLAM and global optimization stages. Subsequently, we assess the performance of the visual-inertial-GNSS fusion on our self-collected Wuhan urban dataset. Monocular camera information is used for all the tests.
VI-A KITTI-360 Dataset